# Polarization of reflected X-ray emission from Sgr A molecular complex: multiple flares, multiple sources?

### v1

@@ Appendix A Systematic uncertainties
The IXPE observations of the Sgr A complex are aimed at faint, spatially and temporarily variable diffuse emission in the vicinity of a number of bright point-like and extended sources. As a result, measurement of the polarization signal and especially mapping of its spatial morphology is very challenging and sensitive to uncertainties in the current knowledge of systematic effects inherent to such observations.
@@
Here we estimate the level of systematic uncertainties that might be present in the data in addition to the purely statistical noise caused by limited count number statistics. In order to do so, we perform a number of jack-knifing experiments by considering variations in the obtained results if various subsets of the full IXPE dataset are included. In particular, one can compare data collected in two epochs separately, by separate Detector Units (DU) of IXPE, as well as spatial variations in excess of the expected statistical noise level. We calculate variance within regions of relevant size (comparable to the brightest in the reflection region) shown in Figures 12 and 13 for Stokes and , respectively.

### v2

@@ Appendix B Systematic uncertainties
The IXPE observations of the Sgr A complex are aimed at faint, spatially and temporally variable diffuse emission in the vicinity of a number of bright point-like and extended sources. As a result, measurement of the polarization signal and especially mapping of its spatial morphology is very challenging and sensitive to uncertainties in the current knowledge of systematic effects inherent to such observations.
@@
Here, we estimate the level of systematic uncertainties that might be present in the data in addition to the purely statistical noise caused by limited count number statistics. In order to do so, we perform a number of jack-knifing experiments by considering variations in the obtained results if various subsets of the full IXPE dataset are included. In particular, one can compare data collected in two epochs separately, by separate Detector Units (DU) of IXPE, as well as spatial variations in excess of the expected statistical noise level. We calculate variance within regions of relevant size (comparable to the brightest in the reflection region) shown in Figs. 12 and 13 for Stokes and , respectively.